OpenStudy (bananas1234):

Refer to triangle ABC. Find the length of side AB. c = 30, b = 17, find a.

OpenStudy (anonymous):

do you know the Pythagorean theorem ?

OpenStudy (bananas1234):

yes\

OpenStudy (anonymous):

you would use that to figure out the side length, do you need help setting it up or anything?

OpenStudy (bananas1234):

yeah i would like some help with that :)

OpenStudy (anonymous):

do you know the equation? for Pythagorean theorem?

OpenStudy (bananas1234):

a^2 + b^2 = c^2

OpenStudy (anonymous):

wait a second, those are angle measurements arent they?

OpenStudy (anonymous):

sorry, I thought they were side lengths,

OpenStudy (bananas1234):

they are side lengths

OpenStudy (anonymous):

oh my goodness, okay, sorry, I confused myself

OpenStudy (bananas1234):

ha ha its ok :)

OpenStudy (anonymous):

let me re draw it, I drew it wrong in the first place

OpenStudy (anonymous):



OpenStudy (anonymous):

so using A2 + B2= C2 (idk how to type the squared thingy) lets plug in the numbers

OpenStudy (bananas1234):

ok so a^2 + 17^2 = 30^2?

OpenStudy (anonymous):

yes, but to make it easier to solve, were gonna flip it so A2= B2+ C2 (because we're trying to find A)

OpenStudy (bananas1234):

ok

OpenStudy (anonymous):

so then we simplify: 17 squared (17*17) = _____

OpenStudy (bananas1234):

289

OpenStudy (anonymous):

I got that too, and then 30 squared (30*30) =___

OpenStudy (bananas1234):

900, sorry if im slow to respond open study is being really laggy for me.

OpenStudy (anonymous):

yeah, I think its because its a live site, and theres tons of people on. But I got 900 too. So then we add the two numbers together

OpenStudy (bananas1234):

1,189

OpenStudy (anonymous):

yep, so then we're left with A2=1189 next, we square root 1189, this will cancel out A2, so all we're left with is A

OpenStudy (bananas1234):

34.5 (rounded)

OpenStudy (anonymous):

yep, you got it

OpenStudy (bananas1234):

awesome!!